Please solve for (a)-(d)below. The data in the accompanying table represent the rate of return of a certain company stock for 11 months, compared witDetermine the P-value for this hypothesis test. P-value = (Round to three decimal places as needed.) State the appropriate co(c) Assuming the residuals are normally distributed, construct a 90% confidence interval for the slope of the true least-squai Rate of Return Х Month Apr-18 May-18 Jun-18 Jul-18 Aug-18 Sept-18 Oct-18 Nov-18 Dec-18 Jan-19 Feb-19 Rates of return of the
Answer #1

S.No х Y (y-7) 1 2 4.23 3.35 3.28 5.09 (x-x)(y-7) 11.7318 16.6022 3 4 -1.78 -3.2 1.29 0.54 2.88 2.69 un 6 7 8 9 3.58 1.48 -4.

a)

ß0 =1.2300

ß1 =0.7648

b)

correlation r='Sxy/(√Sxx*Syy) = 0.663
test stat t= r*(√(n-2)/(1-r2))= 2.659
P value    = 0.026 (from excel:tdist(2.659,9,2)

O A. Reject Ho. There is sufficient evidence to conclude that a linear relation exists between the rate of return of the inde

c)

std error of slope =se(β1) =s/√Sxx= 0.2876
for 95 % CI value of t= 2.262 (from excel:tinv(0.05,9)
margin of error E=t*std error = 0.6506
lower bound=b1+E= 0.1142
Upper bound=b1+E= 1.4154

d)

predicted val=1.23+3.15*0.7648= 3.639